OpenStudy (anonymous):

4. The hypotenuse of a 30°- 60°- 90° right triangle measures 22 cm. What are the lengths of the longer leg and shorter leg? 4. The hypotenuse of a 30°- 60°- 90° right triangle measures 22 cm. What are the lengths of the longer leg and shorter leg? @Mathematics

OpenStudy (anonymous):

sin(30)= opp/hyp = x/22 22*sin(30)=opposite 22*sin(60) = opposite

OpenStudy (anonymous):

22sin(30) = short leg 22sin(60) = long leg

OpenStudy (anonymous):

how did u gwt tht

OpenStudy (anonymous):

let me draw a picture

OpenStudy (anonymous):

ok

OpenStudy (anonymous):